दो संख्याएँ ऐसि है कि उनका अंतर, उनका योग तथा उनका गुणनफल 1:7:24 के अनुपात में है| संख्याओं का गुणनफल है?
Answer by ikleyn(52781)

.
I asked GOOGLE to translate it.

GOOGLE detected "hindi" and created this translation:
    Two numbers Aesi their difference, their sum and their product is in the ratio of 1: 7: 24 | Multiplied by the numbers?

But it is still unclear to me . . .
--------------------


I will suppose the following formulation:
    Find two numbers such that their difference,  their sum and their product are in the ratio of  1 : 7 : 24.

Then  the answer  is:   these numbers are  8 and  6.

Check:  the difference is  D = 8 - 6 = 2 ,

        the sum is         S = 8 + 6 = 14 ,   and

        the product is     P = 8*6 = 48.

        Therefore,  the ratio  D : S : P = 1 : 7 : 24.


Answer.  The numbers are  8  and  6.